OpenStudy (blockcolder):

Given are non-zero non-orthogonal vectors \(\mathbf{A}, \mathbf{B}\). \(\mathbf{B}\) has components parallel and perpendicular to \(\mathbf{A}\) Prove using fundamental ideas and not simply by the manipulation of identities that \(\mathbf{B}\) can be written as: \[\mathbf{B}=\frac{\mathbf{A}(\mathbf{A}\cdot\mathbf{B})}{|\mathbf{A}|^2}+\frac{(\mathbf{A}\times\mathbf{B})\times\mathbf{A}}{|\mathbf{A}|^2}\] I'm having a little bit of trouble deriving the second part. Any hints?

OpenStudy (experimentx):

\[ \frac{(A \times B )\times A}{|A|^2} = -\frac{(A \cdot B)A - (A \cdot A)B}{|A|^2}\]

OpenStudy (blockcolder):

We were told not to use that identity.

OpenStudy (experimentx):

i guess i can't help it

OpenStudy (anonymous):

What do you mean by fundamental ideas?

OpenStudy (blockcolder):

I mean things like the concept of vector addition, subtraction, dot product, cross product, scalar projections, vector projections. Also, the following were allowed: \[\mathbf{a}\cdot\mathbf{b}=|\mathbf{a}||\mathbf{b}|\cos\theta\\ |\mathbf{a}\times\mathbf{b}|=|\mathbf{a}||\mathbf{b}|\sin\theta\] But no other identities are.

OpenStudy (asnaseer):

Maybe this could lead somewhere fruitful - you are told that \(\mathbf{B}\) has some components parallel to \(\mathbf{A}\) and some which are perpendicular to it. So you could write \(\mathbf{B}\) as:\[\mathbf{B}=\alpha\mathbf{A}+\mathbf{C}\]where \(\mathbf{C}\) is perpendicular to \(\mathbf{A}\) such that:\[\mathbf{C}\cdot\mathbf{A}=0\]

OpenStudy (asnaseer):

BTW: \(\alpha\) is just some scalar constant. For example, using this we get:\[\mathbf{A}\cdot\mathbf{B}=\mathbf{A}\cdot(\alpha\mathbf{A}+\mathbf{C})=\alpha\mathbf{A}\cdot\mathbf{A}+\mathbf{A}\cdot\mathbf{C}=\alpha|\mathbf{A}|^2\]

OpenStudy (asnaseer):

\[\therefore\frac{\mathbf{A}\cdot\mathbf{B}}{|\mathbf{A}|^2}=\alpha\]

OpenStudy (asnaseer):

\[\therefore\frac{\mathbf{A}(\mathbf{A}\cdot\mathbf{B})}{|\mathbf{A}|^2}=\alpha\mathbf{A}\]

OpenStudy (asnaseer):

Now you just need to show that:\[\frac{(\mathbf{A}\times\mathbf{B})\times\mathbf{A}}{|\mathbf{A}|^2}=\mathbf{C}\]and you have proved the identity

OpenStudy (blockcolder):

And therein lies my problem. I derived the first part a little differently than you did, but it gives the same result.

OpenStudy (asnaseer):

\[\mathbf{A}\times\mathbf{B}=\mathbf{A}\times(\alpha\mathbf{A}+\mathbf{C})=\alpha\mathbf{A}\times\mathbf{A}+\mathbf{A}\times\mathbf{C}=|\mathbf{A}||\mathbf{C}|\mathbf{y}\]where \(\mathbf{y}\) isa unit vector perpendicular to \(\mathbf{A}\) and \(\mathbf{C}\). Therefore:\[(\mathbf{A}\times\mathbf{B})\times\mathbf{A}=|\mathbf{A}||\mathbf{C}|\mathbf{y}\times\mathbf{A}=|\mathbf{A}||\mathbf{C}||\mathbf{y}||\mathbf{A}|\mathbf{c}=|\mathbf{A}|^2|\mathbf{C}|\mathbf{c}=|\mathbf{A}|^2\mathbf{C}\]therefore:\[\frac{(\mathbf{A}\times\mathbf{B})\times\mathbf{A}}{|\mathbf{A}|^2}=\mathbf{C}\]

OpenStudy (asnaseer):

where \(\mathbf{c}\) is the unit vector in the direction of \(\mathbf{C}\)

OpenStudy (blockcolder):

Wait, I don't understand why \[\mathbf{A}\times\mathbf{C}=|\mathbf{A}||\mathbf{C}|\mathbf{y}\]

OpenStudy (asnaseer):

its because \(\mathbf{A}\) and \(\mathbf{C}\) are perpendicular to one another

OpenStudy (asnaseer):

\[|\mathbf{A}\times\mathbf{C}|=|\mathbf{A}||\mathbf{C}|\sin(90)=|\mathbf{A}||\mathbf{C}|\]

OpenStudy (blockcolder):

Ah, so essentially, the trick is to express \(\mathbf{B}=\alpha\mathbf{A}+\mathbf{C}\) from the get-go. Thanks for the help!

OpenStudy (asnaseer):

yw :)